Use elimination method to solve the linear system Xi + 2x2-4x3-9 4. 2-x3-2
Answer #1

Adding equation (1) and twice the equation (2), we get

x_{1}+10x_{1}+2x_{2}-2x_{2}-4x_{3}+4x_{3}=9+2\Rightarrow 11x_{1}=11\Rightarrow x_{1}=1

equation (2) becomes as 5(1)-x_{2}+2x_{3}=1\Rightarrow -x_{2}+2x_{3}=-4

Adding equation (2) and equation (3), we get

-x_{2}+x_{2}+2x_{3}-x_{3}=-4+2\Rightarrow x_{3}=-2

equation (3) becomes as x_{2}-(-2)=2\Rightarrow x_{2}=2-2\Rightarrow x_{2}=0

Therefore x_{1}=1,x_{2}=0 and x_{3}=-2